OpenStudy (anonymous):

Check my answer please? Find the midpoint of CM C(-5,7) M(-2,1). My answer is; the midpoint is (-3.5,4)

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

nice job

OpenStudy (anonymous):

thank you! :D

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

may I ask, I've forgotten how to find a length segment, which steps to use, for example on this problem I did (-2+-5/2,1+7/2)=(-7/2,8/2)=(-3.5,4) as in (x1+x2/2,y1+y2/2)=(x/2,y/2)=(x,y) if you can help me what's the equation to find a length segment? on this? http://prntscr.com/3mmcix

jimthompson5910 (jim_thompson5910):

are you familiar with the pythagorean theorem?

OpenStudy (anonymous):

yes so far I've figured out (AB)^2=5^2+2^2, (AB)^2=29 but how do I simplify it?

jimthompson5910 (jim_thompson5910):

you want the length of AB so you want to know the value of AB

jimthompson5910 (jim_thompson5910):

which means you need to completely isolate AB

jimthompson5910 (jim_thompson5910):

how do you do that?

OpenStudy (anonymous):

well that's the step I've forgotten

jimthompson5910 (jim_thompson5910):

what operation undoes squaring?

OpenStudy (anonymous):

square root? o.O

jimthompson5910 (jim_thompson5910):

so you apply the square root to both sides

jimthompson5910 (jim_thompson5910):

the square root of AB^2 is just AB (this works since AB > 0)

OpenStudy (anonymous):

I might just sound stupid right now but I did \[\sqrt{29}\] and the answer came up was 5.3 it doesn't seem right to me

jimthompson5910 (jim_thompson5910):

it's more like 5.3851648 which rounds to 5.4 so you have the right idea

OpenStudy (anonymous):

so 5.4 is the right answer?

OpenStudy (anonymous):

my question is the steps I did, was it correct?

jimthompson5910 (jim_thompson5910):

yes, you apply the square root to both sides to get \[\Large AB = \sqrt{29}\]

jimthompson5910 (jim_thompson5910):

then you use a calculator to get this approximation \[\Large AB \approx 5.3851648071345\]

jimthompson5910 (jim_thompson5910):

and you round however you need to

OpenStudy (anonymous):

alrighty! thank you very much once again, you've been a big help to me :D

jimthompson5910 (jim_thompson5910):

you're welcome and I'm glad to be of help